candyplays321:

What is the value of the expression shown below? 7 + (10 − 4)2 ÷ 4 × 1 over 2 to the power of 3 8 8 and 1 over 6 8 and 1 over 8 79

Alexis1415:

Okay so this is a lot of steps so bare with me for a moment.

Alexis1415:

\[7+(10-4)^{2} : 4 \times \frac{ 1 }{ 2^{3}} = 7 + 36 : 4 \ \frac{ 1 }{ 8 }\]

Alexis1415:

Theres more one second

Alexis1415:

\[= 7+9 \times \frac{ 1 }{ 8 } = 7+\frac{ 9 }{ 8 } = 7+ 1\frac{ 1 }{ 8 } = 8\frac{ 1 }{ 8 }\]

candyplays321:

that it?

Alexis1415:

@candyplays321 wrote:
that it?
yeah sorry i was switching classes

candyplays321:

wait what

candyplays321:

8 8 and 1 over 6 8 and 1 over 8 79